solve the system: x^2-144y^2=144
x^2+y^2=144

Answer by venugopalramana(3286)

solve the system: x^2-144y^2=144.....................1
x^2+y^2=144..........................2
eqn.2 - eqn.1
x^2+y^2-x^2+144y^2=144-144=0
145y^2=0
y^2=0
y=0
x^2=144
x=+12 or -12
hence the solution sets are
x=12&y=0
x=-12&y=0